The following data were obtained from a repeated-measures research study. What is the value of MD for these data?

Subject 1st 2nd

#1 10   15

#2 4   8

#3 7    5

#4 6   11​

a.

​4.5

b.

​3

c.

​3.5

d.

​4
Answer #1
S. No 1st 2nd diff:(d)=x1-x2
1 10 15 5
2 4 8 4
3 7 5 -2
4 6 11 5
total = Σd=12
mean dbar= d̅                           = 12/4=3.00

from above MD =3

option b is correct